OpenStudy (wiggler):

Find the horizontal asymptote of (x^2)/((x^4)+2)

OpenStudy (wiggler):

\[\frac{ x^2 }{ \sqrt{x^4+2} }\]

OpenStudy (anonymous):

hint: x^2 = sqrt(x^4)

OpenStudy (anonymous):

If the degree of the denominator is bigger than the degree of the numerator, the horizontal asymptote is the x-axis (y=0)

OpenStudy (wiggler):

how do i figure out if \[\sqrt{x^4+2}\] is a bigger or smaller degree?

OpenStudy (anonymous):

What is the degree in the denominator?

OpenStudy (wiggler):

4?

OpenStudy (wiggler):

Or is it 2, since it's sqrt?

OpenStudy (anonymous):

correct, it's 4

OpenStudy (anonymous):

do you understand what horizontal asymptote is?

OpenStudy (wiggler):

Yep, the limit as x tends to - or + infinity (where y get close to in the distance)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so here are some steps to help you find the horizontal asymptote: 1) Put equation or function in y= form. 2) Multiply out (expand) any factored polynomials in the numerator or denominator. 3) Remove everything except the terms with the biggest exponents of x found in the numerator and denominator. These are the "dominant" terms.

OpenStudy (wiggler):

I'm just having trouble expanding the denominator \[\sqrt{x^4+2}\]

OpenStudy (anonymous):

Refer to the attachment.